A body is projected at an angle 45 degrees with the horizontal with velocity 40 root2 m/s. Its resulatant displacement after 2secs isa.80m b.100m c.640m d.60m

Grade:12

1 Answers

Adarsh
6 years ago
We can calculate it by finding horizontal and vertical displacement of particles .Horizontal and vertical component of velocity is equal to 40m/s.Vertical Displacement of particle after 2 second = 40.2-1/2.10.2.2=60mHorizontal displacement of particle after 2 second = 40.2-0= 80mNow net displacement of particle = √ (80)^2 + (60)^2 = 100m.Hence (b) is the correct option.